# Detection of Aerial Spoofing Attacks to LEO Satellite Systems via Deep Learning

### v1

@@ V Satellite Spoofing Detection
TABLE II: Notation and description.
V-A Image Generation
@@
Our solution analyzes raw IQ samples, collected through a device capable of extracting PHY-layer data from the received signal, e.g., a SDR or a spectrum analyzer. We divide the stream of IQ samples into subsets (non-overlapping chunks) of $N$ samples, i.e., $\{(I_{k},Q_{k})\}_{k=1}^{N}$ (see Sec. VII-B for an evaluation of the impact of $N$ on the performance of our solution), and visualize them through an IQ constellation diagram, with the I component on the x-axis and the Q component on the y-axis. Then we divide the diagram into a $P\times Q$ grid, with $P$ and $Q$ chosen according to the methodology used for spoofing detection (see Sec. V-B). We count how many samples fall into each tile of the grid, so generating a histogram $H$. Denoting with $\mathbf{b_{I}}$ and $\mathbf{b_{Q}}$ the set of edges of the tiles of the histogram along X and Y, respectively, the value of the histogram tile $H_{i,j}$ follows Eq. 2.
@@
V-B Spoofing Detection via Autoencoders
We identify a spoofing attack using sparse AEs, by detecting anomalies in the received RF signal due to the different fading phenomena affecting the communication link (satellite vs aerial). In particular, the shape of the IQ clouds at the receiver when considering benign and malicious signals is different— the signal received from the spoofer installed on the drone has more statistical variance around the expected values than the signal received from the satellites. We attribute this phenomenon to the different fading phenomena affecting the received signals, as the signal generated by the satellites comes from a different altitude than the one originated by the spoofer on the drone, being subject to different fading phenomena. Figure 3 shows the architecture of the adopted AE.
@@ V-B Spoofing Detection via Autoencoders
We provide images (generated through the procedure described in Sec. V-A) as input to an encoder using a logarithmic sigmoid activation function with $L=16$ neurons. Such an operation generates a compressed latent representation of the input image, characterized by $L=16$ dimensions. We feed such latent representation vectors to a decoder using a linear decoder transfer function, using $S=50,176$ neurons. Our AE architecture uses two hidden layers and the sparsity regularization technique. We obtain as output an image of the same size as the input ($P\times Q$), which we convert into a matrix and compare with the input image through the MSE loss function, as in Eq. 1 (Sec. II). Our approach involves training and testing. For the training process, we use only images generated from IQ samples captured from legitimate LEO satellite communications, generating a set of MSE values characterizing the regular pattern of the communication link, namely $\mathbf{MSE_{train}}$. We use such values to compute the threshold $\tau$, used to differentiate images generated from legitimate satellites from spoofed ones. An image with $MSE\leq\tau$ is predicted to be generated from legitimate transmitters (LEO satellites), while an image with $MSE>\tau$ is predicted to be generated from unauthorized entities (adversary). For computing $\tau$, we use Eq. 4, as proposed by the authors in [40] (Sec.4.1), so as to take into account as much as possible the variability of legitimate communication patterns due to noise.
@@
Note that we empirically select a hidden size value of $16$, a sparsity regularization term of $0.5$, an L2-regularization term of $0.01$, and $250$ epochs for the training.
VI Data Collection
In this section, we describe the experimental setup used for data collection (Sec. VI-A), the experiment settings (Sec. VI-B), and we report on a preliminary statistical analysis of our measurements (Sec. VI-C).
VI-A Experimental Setup
@@ VI-B Experiment Settings
We believe that such scenarios cover the most realistic cases of aerial spoofing attacks to LEO satellite receivers deployed outdoors, making our results generalizable to any receivers deployed outdoors in a scenario like the one shown in Fig. 6. Moreover, our experiments cover multiple altitudes and fading dynamics, addressing possible concerns about the limited scalability and robustness of our experimental findings. Note that, to avoid interfering with other potential IRIDIUM receivers in the area, we selected a TX power gain value on the transmitter low enough not to generate interference further away than $500$ m, while being high enough to still receive messages from the maximum altitude of $120$ m. Finally, note that we did not perform new measurements for ground-based spoofing attacks, as such scenarios are already covered by the literature [10].
VI-C Data Analysis
Our collected dataset includes more than $10,000$ legitimate IRIDIUM messages. Moreover, we collected more than $1,000$ spoofed IRIDIUM messages for each scenario and tested altitude, as summarized in Tab. III, for a total number of approx. $24,000$ messages. Our dataset is available as open source at [12].
@@ VI-C Data Analysis
TABLE III: Collected spoofed messages.
@@
VII Experimental Performance Analysis
In this section, we report the main results of our experimental performance assessment. We explain the metrics used for our results in Sec. VII-A and report our results in Sec. VII-B.
VII-A Metrics
In the following, we report the performance of all the tested solutions using the Area Under-the-Curve (AUC) of the Receiver Operating Characteristic (ROC) curve. This metric ranges between the values 00 and $1$, with $0.5$ representing a random guess and $1$ indicating a perfect classification. Note that the ROC AUC measures the quality of a classifier independently of the threshold selection process, and it is recommended when the datasets are unbalanced, as it evaluates the trade-off between sensitivity (true positive rate) and specificity (true negative rate) in various threshold settings [46][47]. For each reported figure, we train our proposed solution as outlined in Sec. V, using $80$% of the images generated from legitimate samples. We validate the model using the remaining $20$% and then test it using spoofed data acquired in a particular scenario, as described in Sec. VI. We use K-fold validation with $K=5$ for each configuration and report the quantiles 0.05, 0.5, and 0.95 for the AUC values. Finally, we compare our proposed solution against two benchmarks, i.e., the solution in [10], using a binary Fully Convolutional Data Descriptor (FCDD) to detect spoofing attacks [48], and the proposal in [8], using the neural network Resnet-18 to authenticate satellite wireless messages. Note that the proposal in [10] considers the same problem tackled in this manuscript. Thus, no changes are needed to ensure a fair comparison. Conversely, to allow for a fair comparison between our setup and the proposal in [8], we use the neural network classifier Resnet-18 to distinguish between satellite (authorized) and drone (unauthorized) transmissions using binary classification.
@@ VII-A Metrics
VII-B Results
We report in Fig. 10 (a-c) the performance of our solution in the three investigated scenarios, using various numbers of samples per image $N$ (recall Sec. V-A).
@@ VII-B Results
We further consider the configuration of our solution using $N=250$ samples per image (often performing the worst in Fig. 10), and we also compare the performance of our proposed solution to the benchmarks proposed by Oligeri et al. in [10] and [8]. We report the results of our comparison in Figs. 11(a-c).
@@
VIII Conclusion and Future Work
@@
In this paper, we proposed a new solution for detecting aerial spoofing attacks to LEO satellite communication systems, using a sparse Autoencoder. Our proposed solution requires converting received Physical information (IQ samples) into images and then using an Autoencoder to identify which of them come from the spoofer. We validated our solution using data collected through an extensive real-world measurement campaign, encompassing actual spoofing attacks of a reference LEO satellite system (IRIDIUM) via a Software-Defined Radio installed on a drone and flown in a remote area via several attack scenarios (static, transmitter movement, transmitter and receiver movement). Our proposed solution can detect aerial spoofing attacks with a minimum ROC AUC value of $0.75$, systematically outperforming competing approaches. We also released the dataset collected for this study as open source to the research community, contributing to stimulating further real-world applied research on satellite security. Note that, in our work, we could not fly the drone beyond the maximum altitude of $120$ m imposed by EU regulations. Also, we did not carry out tests in an urban scenario, where the presence of man-made construction and people’ movement might change the profile of the signal received from satellites. In line with such limitations, in future work, we plan to test our solution in urban environments, so to assess the impact of man-made constructions on the effectiveness of our approach. Moreover, we will work on an analytical channel model for LEO satellite transmissions to analyze the performance of our solution in the presence of attackers circumventing law regulations on the maximum allowed flight altitude of drones.
Acknowledgements
This work has been partially supported by the INTERSECT project, Grant No. NWA.1162.18.301, funded by the Netherlands Organization for Scientific Research (NWO). Any opinions, findings, conclusions, or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of NWO. Moreover, this publication was made possible by the NPRP12C-0814-190012-SP165 awards from the Qatar National Research Fund (a member of Qatar Foundation).

### v2

@@ 5 Satellite Spoofing Detection
Table 2: Notation and description.
5.1 Image Generation
@@
Our solution analyzes raw IQ samples, collected through a device capable of extracting Physical (PHY)-layer data from the received signal, e.g., a SDR or a spectrum analyzer. We divide the stream of IQ samples into subsets (non-overlapping chunks) of $N$ samples, i.e., $\{(I_{k},Q_{k})\}_{k=1}^{N}$ (see Sec. 7.2 for an evaluation of the impact of $N$ on the performance of our solution), and visualize them through an IQ constellation diagram, with the I component on the x-axis and the Q component on the y-axis. Then we divide the diagram into a $P\times Q$ grid, with $P$ and $Q$ chosen according to the methodology used for spoofing detection (see Sec. 5.2). We count how many samples fall into each tile of the grid, so generating a histogram $H$. Denoting with $\mathbf{b_{I}}$ and $\mathbf{b_{Q}}$ the set of edges of the tiles of the histogram along X and Y, respectively, the value of the histogram tile $H_{i,j}$ follows Eq. 2.
@@
5.2 Spoofing Detection via Autoencoders
We identify a spoofing attack using sparse AEs, by detecting anomalies in the images generated from the received RF signal due to different fading phenomena affecting the communication link (satellite vs aerial). In particular, the shape of the IQ clouds at the receiver when considering benign and malicious signals is different— the signal received from the spoofer installed on the drone has more statistical variance around the expected values than the signal received from the satellites. We attribute this phenomenon to the different fading phenomena affecting the received signals, as the signal generated by the satellites comes from a different altitude than the one originated by the spoofer on the drone, being subject to different fading.
According to such considerations, the key design objective of our solution is to learn an efficient, compressed representation (encoding) of the input IQ data, for the purpose of feature learning, while being able to reconstruct the original input as accurately as possible from this encoding. Using mathematical notation, considering: (i) the input $\mathbf{x}$ (image generated from $N$ IQ samples), (ii) the encoder function $f_{\theta}$, with $\theta$ being a parameter that maps $\mathbf{x}$ over the latent representation $\mathbf{z}$, (iii) the decoder function $g(\phi)$, being $\phi$ a parameter that is used for decoding $\mathbf{z}$ into a reconstruction of the input $\mathbf{\tilde{x}}$, and (iv) $\hat{\mathbf{x}}=g_{\phi}(f_{\theta}(\mathbf{x}))$ the reconstruction of the input, the objective of our autoencoder is to minimize the reconstruction loss as defined in [43] over the dataset of the training images $D$, according to Eq. 4:
$\min_{\theta,\phi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\left[\mathcal{L}(\mathbf{x},\hat{\mathbf{x}})\right]=\min_{\theta,\phi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\left[\mathcal{L}(\mathbf{x},g_{\phi}(f_{\theta}(\mathbf{x})))\right]\|_{2}^{2},$ (4)
where $\mathcal{L}$ is the reconstruction loss function introduced in Eq. 1.
Figure 3 shows the architecture of the adopted AE.
@@ 5.2 Spoofing Detection via Autoencoders
We provide images (generated through the procedure described in Sec. 5.1) as input to an encoder using a logarithmic sigmoid activation function with $L=16$ neurons. Such an operation generates a compressed latent representation of the input image, characterized by $L=16$ dimensions. Specifically, our encoder generates the latent representation $\mathbf{z}$ according to Eq. 5:
$\mathbf{z}=\sigma\left(W^{(1)}\mathbf{x}+\mathbf{b}^{(1)}\right),$ (5)
where $(W^{(1)}\in\mathbb{R}^{16\times n})$ is the encoder weight matrix, $(\mathbf{b}^{(1)}\in\mathbb{R}^{16})$ is the bias vector of the encoder and $\sigma(s)=\frac{1}{1+e^{-s}}$ is the sigmoid activation function. We feed such a latent representation vector $\mathbf{z}$ to a decoder using a linear decoder transfer function, using $S=50,176$ neurons. Our AE architecture uses two hidden layers, the optimizer algorithm Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm, and the sparsity regularization technique. Thus, the operations performed by the decoder to obtain the reconstructed input $\hat{\mathbf{x}}$ can be expressed as in Eq. 6.
$\mathbf{z}^{(2)}$ $=\phi^{(2)}\left(W^{(2)}\mathbf{z}^{(1)}+\mathbf{b}^{(2)}\right),\quad\mathbf{z}^{(2)}\in\mathbb{R}^{h_{2}}$ (6)
$\mathbf{z}^{(3)}$ $=\phi^{(3)}\left(W^{(3)}\mathbf{z}^{(2)}+\mathbf{b}^{(3)}\right),\quad\mathbf{z}^{(3)}\in\mathbb{R}^{h_{3}}$
$\hat{\mathbf{x}}$ $=W^{(4)}\mathbf{z}^{(3)}+\mathbf{b}^{(4)},\quad\hat{\mathbf{x}}\in\mathbb{R}^{50176}$
where $(\phi^{(2)},\phi^{(3)})$ are activation functions for decoder hidden layers, $(W^{(2)},W^{(3)},W^{(4)})$ are weights for each decoder layer, $(\mathbf{b}^{(2)},\mathbf{b}^{(3)})$, are biases for each decoder layer, and $\mathbf{z}^{(2)}$ and $\mathbf{z}^{(3)}$ represent the intermediate activations (or outputs) of each decoder hidden layer, namely the hidden states. Note that, in line with general principles of autoencoders (see [43]), for a deep network with $L$ layers there are $L+1$ weight matrices, i.e., one for each layer and one final weight matrix to project from the last hidden layer to the output. We obtain as output $\hat{\mathbf{x}}$ an image of the same size as the input ($P\times Q$), which we convert into a matrix and compare with the input image through the MSE loss function, as in Eq. 1 (Sec. 2).
Our approach involves training and testing. For the training process, we use only images generated from IQ samples captured from legitimate LEO satellite communications, generating a set of MSE values characterizing the regular pattern of the communication link, namely $\mathbf{MSE_{train}}$. We use such values to compute the threshold $\tau$, used to differentiate images generated from legitimate satellites from spoofed ones. An image with $MSE\leq\tau$ is predicted to be generated from legitimate transmitters (LEO satellites), while an image with $MSE>\tau$ is predicted to be generated from unauthorized entities (adversary). For computing $\tau$, we use Eq. 7, as proposed by the authors in [44] (Sec.4.1), so as to take into account as much as possible the variability of legitimate communication patterns due to noise.
@@
Note that we empirically select a hidden size value of $16$, a sparsity regularization term of $0.5$, an L2-regularization term of $0.01$, and $250$ epochs for the training. Such values were selected through experimental fine-tuning based on the available data (see later section).
6 Data Collection
In this section, we describe the experimental setup used for data collection (Sec. 6.1), the experiment settings (Sec. 6.2), and we report on a preliminary statistical analysis of our measurements (Sec. 6.3).
6.1 Experimental Setup
@@ 6.2 Experiment Settings
We believe that such scenarios cover the most realistic cases of aerial spoofing attacks to LEO satellite receivers deployed outdoors, making our results generalizable to any receivers deployed outdoors in a scenario like the one shown in Fig. 6. Moreover, our experiments consider various movement patterns, accounting for heterogeneous noise patterns, multipath, and interference conditions characterizing real-world satellite communication channels. Our experiments also cover multiple altitudes and fading dynamics, addressing possible concerns about the limited scalability and robustness of our experimental findings. Note that, to avoid interfering with other potential IRIDIUM receivers in the area, we selected a TX power gain value on the transmitter low enough not to generate interference further away than $500$ m, while being high enough to still receive messages from the maximum altitude of $120$ m. Finally, note that we did not perform new measurements for ground-based spoofing attacks, as such scenarios are already covered by the literature [11].
6.3 Data Analysis
Our collected dataset includes more than $10,000$ legitimate IRIDIUM messages. Moreover, we collected more than $1,000$ spoofed IRIDIUM messages for each scenario and tested altitude, as summarized in Tab. 3, for a total number of approx. $24,000$ messages. Our dataset is available as open source at [13].
@@ 6.3 Data Analysis
Table 3: Collected spoofed messages.
@@
7 Experimental Performance Analysis
In this section, we report the main results of our experimental performance assessment. We explain the metrics used for our results in Sec. 7.1 and report our results in Sec. 7.2.
7.1 Metrics
In the following, we report the performance of all the tested solutions using the Area Under-the-Curve (AUC) of the Receiver Operating Characteristic (ROC) curve. The ROC curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at various threshold settings. Therefore, the ROC curve inherently depends on how false positives (FP) vary with different classification thresholds. The AUC is the area under this ROC curve, so it summarizes how well the classifier distinguishes between classes across all thresholds, taking into account both true positives and false positives [50]. This metric ranges between the values 00 and $1$, with $0.5$ representing a random guess and $1$ indicating a perfect classification. Note that the ROC AUC measures the quality of a classifier independently of the threshold selection process, and it is recommended when the datasets are unbalanced, as it evaluates the trade-off between sensitivity (true positive rate) and specificity (true negative rate) in various threshold settings [51][52].
For each reported figure, we train our proposed solution as outlined in Sec. 5, using $80$% of the images generated from legitimate samples. We validate the model using the remaining $20$% and then test it using spoofed data acquired in a particular scenario, as described in Sec. 6. We use K-fold validation with $K=5$ for each configuration and report the quantiles 0.05, 0.5, and 0.95 for the ROC AUC values. Finally, we compare our proposed solution against two benchmarks, i.e., the solution in [11], using a binary Fully Convolutional Data Descriptor (FCDD) to detect spoofing attacks [53], and the proposal in [9], using the neural network Resnet-18 to authenticate satellite wireless messages. Note that the proposal in [11] considers the same problem tackled in this manuscript. Thus, no changes are needed to ensure a fair comparison. Conversely, to allow for a fair comparison between our setup and the proposal in [9], we use the neural network classifier Resnet-18 to distinguish between satellite (authorized) and drone (unauthorized) transmissions using binary classification.
@@ 7.1 Metrics
7.2 Results
We report in Fig. 10 (a-c) the performance of our solution in the three investigated scenarios, using various numbers of samples per image $N$ (recall Sec. 5.1).
@@ 7.2 Results
We also measured the overhead of our solution, i.e., the time required to generate an image from incoming IQ samples, namely $\tau_{img}$, and the time to classify such image as benign or spoofed, namely $\tau_{ae}$. We also estimated the memory requirements of our solution, i.e., the size on disk of the model ($M_{ae}$) and the size of the image on disk with various no. of samples $M_{image}$. We report the results in Tab. 4.
Samples per Image (N) $250$ $500$ $1000$
$\tau_{img}$ (ms) 0.0085 0.0089 0.0102
$\tau_{ae}$ (ms) 0.0027 0.0028 0.0032
$M_{ae}$ (MB) 13 13 13
$M_{image}$ (KB) 3.6 4.35 5.76
Table 4: Overhead components of our solution.
We notice that collecting samples necessary to generate an image with $N=250$, $N=500$, or $N=1,000$ samples is in the order of a few micro-seconds ($10$ $\mu$sec with $N=1,000$ samples). Even in the worst case where we use $N=1,000$, the time to classify such an image as benign or malicious on our computing equipment is very limited, i.e., $0.032$ msec, and the generated model requires only $13$ MB of storage space. Therefore, the computational cost of the solution is very limited, as well as its storage requirements are limited for regular computing units. We further consider the configuration of our solution using $N=250$ samples per image (often performing the worst in Fig. 10), and we also compare the performance of our proposed solution to the benchmarks proposed by Oligeri et al. in [11] and [9]. We report the results of our comparison in Figs. 11(a-c).
@@
8 Discussion
Implications for Practice. The results provided in Sec. 7.2 indicate that, through our solution, we can detect spoofing attacks to LEO satellite communication systems originating by drones by analyzing the physical-layer signals at the satellite receiver, and looking for anomalies in the received signal. At the same time, especially when considering $N=50$ samples per image, the average optimal ROC AUC values sometimes do not exceed the value 0.75. While such performances are considerably higher than the state of the art (see Fig. 11), they are not high enough to provide reliable detection in real-world systems. Thus, we foresee that our solution could be used to provide an additional layer of security to existing security systems, rather than as a standalone solution. This is common to many security systems analyzing wireless signals, e.g., Radio Frequency Fingerprinting [9]. We also notice that detection accuracy improves when there are more samples available per message and when the receiver is mobile (see Fig. 11 (c)). Thus, whenever possible, receivers should oversample the incoming signal and move periodically to maximize detection accuracy and reduce false positives.
Limitations. The results reported in Fig. 10 and Fig. 11 suggest that the attacker can increase the chances of bypassing our solution by flying higher, so better mimicking the satellite communication channel. However, due to the regulations in place in our region, in this research, we conducted tests with a drone flying at the maximum altitude of $120$ m—being the maximum allowed altitude [48]. We acknowledge that attackers could neglect such regulations and execute advanced spoofing strategies by flying the drone at a higher altitude. Future research could test our solution with data collected in regions with more relaxed regulations, investigating further the capabilities of advanced attackers. Our tests were also executed in large rural areas, so to minimize safety concerns. Future research could test our solution with data collected in urban environments, to check if reflections caused by man-made constructions could help the attacker, or strengthen the defense strategy.
9 Conclusion and Future Work
@@
In this paper, we proposed a new solution for detecting aerial spoofing attacks to LEO satellite communication systems, using a sparse Autoencoder. Our proposed solution requires converting received Physical information (IQ samples) into images and then using an Autoencoder to identify which of them come from the spoofer. We validated our solution using data collected through an extensive real-world measurement campaign, encompassing actual spoofing attacks of a reference LEO satellite system (IRIDIUM) via a Software-Defined Radio installed on a drone and flown in a remote area via several attack scenarios (static, transmitter movement, transmitter and receiver movement). Our proposed solution can detect aerial spoofing attacks with a minimum ROC AUC value of $0.75$, systematically outperforming competing approaches. We also released the dataset collected for this study as open source to the research community, contributing to stimulating further real-world applied research on satellite security. Note that, in our work, we could not fly the drone beyond the maximum altitude of $120$ m imposed by EU regulations. Also, we did not carry out tests in an urban scenario, where the presence of man-made construction and people’ movement might change the profile of the signal received from satellites. In line with such limitations, in future work, we plan to test our solution in urban environments, so to assess the impact of man-made constructions on the effectiveness of our approach. Moreover, we will work on an analytical channel model for LEO satellite transmissions to analyze the performance of our solution in the presence of attackers circumventing law regulations on the maximum allowed flight altitude of drones. Finally, we will investigate the feasibility of detecting advanced drone-based LEO spoofing attacks onboard other drones and LEO satellites.